OpenStudy (anonymous):

Natasha rides her bike (at a constant speed) for 5 hours, helped by the wind of 3 miles per hour. Pedaling at the same rate, the trip back against the wind takes 11 hours. Find the total round trip distance she traveled. (I have the answer im just not sure how to do the work)

OpenStudy (ddcamp):

Let p be the speed she pedals at and d be the distance We know from the definition of speed that s = d/t → t = d/s Now we have two equations, one for heading out and one for coming back: 5 = d/(p+3) 11 = d/(p-3) Can you solve that system?

OpenStudy (anonymous):

Is that distance over speed plus 3